OpenStudy (anonymous):

If a 30-foot silo casts a 30-foot shadow, what is the measure from the tip of the shadow to the top of the silo?

OpenStudy (anonymous):

30 SQUARED + 30 SQUARED EQUALS H SQUAred. pythagorean theorm.

OpenStudy (anonymous):

Use pythagoras theorem: \[\large a^2 + b^2 = c^2\] \[\large c^2 = 30^2 + 30^2\] Find c here..

OpenStudy (anonymous):

3600

OpenStudy (anonymous):

No..

OpenStudy (anonymous):

wattt 30+30=60 60*60=3600

OpenStudy (anonymous):

30 x 30 equals ?

OpenStudy (anonymous):

\(30^2 + 30^2 = ??\)

OpenStudy (anonymous):

900

OpenStudy (anonymous):

\[\large 30 \times 30 + 30 \times 30 = ??\]

OpenStudy (anonymous):

1800

OpenStudy (anonymous):

Take its square root now..

OpenStudy (anonymous):

42

OpenStudy (anonymous):

Yeah approx..

OpenStudy (anonymous):

so thats the answer

OpenStudy (anonymous):

Yes it is..

OpenStudy (anonymous):

wowwww

OpenStudy (anonymous):

Wooowwwwwww..